Suppose 76% of internet searches used Google. If we take a sample of 4 searches, what is the probability that none of them use Google? Round your answer to 3 decimal places.
Answer #1

Given that ,

p = 0.76

1 - p = 1 - 0.76= 0.24

n =4

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =0 ) = (4 C 0) * 0.760 * (0.24)4

probability =0.003